OpenStudy (anonymous):

How to write >>>y=3x^2-12x+7<<< in vertex form?

OpenStudy (anonymous):

y=(3x^2-12x+___)+7 let's complete the square y=3(x^2-4x+4)+7-12 y=3(x-2)^2-5

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no problem